# ReVision: A Dataset and Baseline VLM for Privacy-Preserving Task-Oriented Visual Instruction Rewriting

### v1

@@ 5 Evaluation Metrics
To evaluate our models in Visual Instruction Rewriting, we use standard NLG metrics (BLEU, METEOR, ROUGE) Sharma et al. (2017) alongside task-specific semantic parsing evaluations. While NLG metrics assess linguistic similarity, they do not capture functional quality in downstream AI systems. Effective rewriting must ensure instructions remain interpretable by semantic parsers extracting intent and arguments Louvan and Magnini (2020). In the absence of a baseline parser for our ontology (Figure 5), we use GPT-4 as a proxy parser to classify intents and extract arguments, simulating an on-device parser. We compare GPT-4-generated parses for reference and model rewrites to verify preservation of intent and structured arguments. We measure the following metrics:
@@ 6 Results and Discussion
Table 3 presents the evaluation results for both baseline models (BL) and our proposed ReVision models across Language Generation (NLG) metrics (ROUGE, BLEU, METEOR) and semantic parsing performance (Intent Accuracy and Argument Mean Jaccard Similarity). We also provide anecdotal examples in Figure 7 in the Appendix to illustrate the strengths and limitations of various models.
As we can see, the baseline models struggle significantly in the rewriting task, not because they are inherently weak, but because they are not explicitly tuned for rewriting. While PaliGemma2-10B and QwenVL-7B have demonstrated strong performance in various vision-language tasks, they are not optimized to follow meta-instructions like rewriting. This is evident in their vanilla versions (BL1a, BL1b), where ROUGE-1 scores remain low (3.4% and 43.7%), BLEU is nearly negligible (0.03% and 12.3%), and Intent Accuracy is poor (16.2% and 50.3%). A key issue is that these models misinterpret the task, often either responding to the instruction directly or attempting to autocomplete the instruction instead of rewriting it. Since many input instructions are imperative, task-oriented instructions, they frequently refuse to generate a rewrite (e.g., replying "I can’t help with that") or incorrectly complete the instruction, significantly degrading both NLG and parsing metrics. Moreover, since they are relatively small models (<10B parameters), they lack the necessary world knowledge and instruction-following capabilities to recognize and execute rewriting as a structured transformation task effectively. Additionally, prompting these models with Self_Caption+EasyOCR metadata (BL2a, BL2b) helps slightly, especially for QwenVL-7B, where Intent Accuracy improves from 50.3% to 61.2%. However, the models still struggle with ROUGE and BLEU scores, reinforcing that generic VLMs require dedicated instruction tuning to handle the rewriting task effectively.
In contrast, our proposed ReVision models, specifically trained for rewriting, significantly outperform all baselines. Even without metadata enhancements, ReVision-BL already outperforms the input-augmented baseline models, achieving ROUGE-1 of 56.9%, BLEU of 27.7%, and Intent Accuracy of 56.5%. This confirms that explicit tuning for rewriting is essential and that even a compact, instruction-tuned VLM can surpass larger models that lack task-specific optimization. These observations are also corroborated by the intent category-wise F1 scores reported in Figure 4.
With metadata, the performance further improves. ReVision-Metadata with GPT4-derived captions and OCR Text achieves 72.4% ROUGE-1, 49.9% BLEU, and 62.4% Intent Accuracy, showing that supplementing the input with extracted text helps models disambiguate multimodal instructions and produce more accurate rewrites. The best-performing model, ReVision-MetadataSelf_Caption+EasyOCR, achieves the highest scores across all metrics, confirming that even lightweight OCR and captioning models can be leveraged to improve rewriting quality. Lastly, the 8-bit quantized version of the best-performing model offers competitive performance to its full-precision counterpart, with only a minor drop in Intent Accuracy (67.6% vs. 71.5%) but a slight improvement in Argument Similarity (79.5%). This demonstrates that 8-bit models can be effectively deployed on resource-constrained devices.
@@
Despite the strong performance of our ReVision model variants, certain limitations prevent even higher accuracy. One major challenge is the loss of fine-grained text details due to image downsampling to $256\times 256$ resolution, making it difficult for the model to capture small but critical information, such as ingredient lists or nutritional facts on product packaging. Additionally, the lack of explicit reference localization in the dataset restricts the model’s ability to map user intent to specific image regions, leading to errors in object disambiguation and instruction alignment. To address these challenges, future work can incorporate bounding box annotations to provide spatial grounding cues for better reference resolution. Processing localized image regions instead of entire downsampled images could reduce information loss, especially for text-heavy visual instructions. This aligns with Pali-Gemma’s short-resolution increase technique Beyer et al. (2024), which enhances fine-grained visual understanding. Despite these limitations, our results confirm that task-specific instruction tuning and metadata augmentation significantly enhance multimodal rewriting, ensuring scalable and efficient on-device deployment.
@@ 7 Conclusion and Future Work
In this work, we explored Visual Instruction Rewriting as a lightweight, privacy-preserving approach to multimodal interaction on AR, VR, and smartphone devices. With a strong emphasis on dataset development, we present a diverse collection of 39,000+ examples covering 14 domains, enabling robust training for on-device instruction rewriting. Our approach ensures that text-only inference is more secure in privacy-sensitive settings by eliminating the need to send personal vision-related images to the server, reducing data exposure risks. Additionally, rewriting removes the necessity of storing images, making multimodal AI systems more efficient and privacy-focused. Our experimental results show that even an 8-bit quantized model maintains strong performance while significantly reducing memory requirements. For future work, we aim to expand data coverage by incorporating more diverse real-world multimodal instructions and introducing multilingual support to enhance accessibility. Furthermore, improving deixis resolution with bounding box annotations and localized image region training will enhance reference grounding while integrating gaze tracking and tactile input can further refine contextual understanding in on-device AI assistants.
@@
While our approach demonstrates strong performance in Visual Instruction Rewriting, several limitations remain. First, image downsampling to $256\times 256$ resolution can lead to the loss of fine-grained text details, affecting instructions that rely on small-font information, such as nutritional labels or product specifications. Second, deictic reference resolution remains challenging, especially in images with multiple similar objects where the model lacks explicit localization cues. The absence of bounding box annotations in our dataset limits the model’s ability to disambiguate references, leading to errors in object-grounded instructions. Additionally, while our model is lightweight and optimized for on-device execution, it still lags behind larger VLMs in handling complex multimodal instructions requiring deep reasoning and external world knowledge. Lastly, our dataset, while diverse across 14 domains, is monolingual, limiting applicability to multilingual and culturally varied settings. Future work can address these challenges by increasing dataset coverage, incorporating localized image region processing, and adding bounding box annotations to improve reference resolution and multimodal grounding.
Ethics Statement

### v2

@@ 5 Evaluation Metrics
To evaluate our models in Visual Instruction Rewriting, we use standard NLG metrics (BLEU, METEOR, ROUGE) sharma2017nlgeval alongside task-specific semantic parsing evaluations. While NLG metrics assess linguistic similarity, they do not capture functional quality in downstream AI systems. Effective rewriting must also ensure instructions remain interpretable by semantic parsers extracting intent and arguments louvan2020recent. In the absence of an existing parser tailored to our ontology (Figure 5), we employ GPT-4 as a proxy to simulate an on-device parser for intent classification and argument extraction. To evaluate intent and structure preservation, we compare GPT-4-generated parses for both reference and model-generated rewrites. For clarity, we present a collapsed view of intents and arguments. The following metrics are used for evaluation
@@ 6 Results and Discussion
Table 3 presents the evaluation results for both baseline models (BL) and our proposed ReVision models across Language Generation (NLG) metrics (ROUGE, BLEU, METEOR) and semantic parsing performance (Intent Accuracy and Argument Mean Jaccard Similarity). We also provide anecdotal examples in Figure 7 to illustrate the strengths and limitations of various models.
Both TextOnly baseline variants performed significantly worse than ReVision, highlighting the value of multimodal inputs. These models struggled with proper nouns and named entities from captions and OCR, showing high sensitivity to metadata quality. Midsize baseline VLMs underperformed not due to weak modeling but due to lack of tuning for rewriting. Though PaliGemma2-10B and QwenVL-7B perform well on vision-language tasks, they are not optimized for meta-instruction following, as seen in their vanilla versions (BL1a, BL1b) with low ROUGE-1 (3.4%, 43.7%), negligible BLEU (0.03%, 12.3%), and poor Intent Accuracy (16.2%, 50.3%). They often misinterpret rewriting as direct response or autocompletion, especially with imperative inputs, leading to refusal (“I can’t help with that”) or incorrect completions—hurting NLG and parsing metrics. Their small size (<10B parameters) limits instruction-following and world knowledge needed for structured rewriting. Adding Self_Caption+EasyOCR metadata (BL2a, BL2b) slightly helps, notably for QwenVL-7B (Intent Accuracy: 50.3% $\rightarrow$ 61.2%), but ROUGE and BLEU remain low, showing the need for instruction tuning. The fine-tuned SmolVLM-256M baseline also underperforms with default tuning, often over-explaining by adding unnecessary descriptions and artifacts, likely due to pretraining on detailed tasks (e.g., video narration). While suboptimal here, SmolVLM remains a promising candidate with targeted tuning and prompting.
In contrast, our proposed ReVision models, explicitly trained for rewriting, substantially outperform all baselines, demonstrating the importance of task-specific tuning. Even without metadata, ReVision-BL exceeds input-augmented baselines with ROUGE-1 of 56.9%, BLEU of 27.7%, and Intent Accuracy of 56.5%, highlighting that a compact, instruction-tuned VLM can surpass larger, non-specialized models—an observation further supported by the intent category-wise F1 scores in Figure 4. Incorporating metadata yields additional gains: ReVision-Metadata, enhanced with GPT-4-derived captions and OCR text, achieves 72.4% ROUGE-1, 49.9% BLEU, and 62.4% Intent Accuracy, confirming that extracted text aids in resolving multimodal ambiguities. The top-performing model, ReVision-Metadata-Self_Caption+EasyOCR, achieves the highest scores across all metrics, showing that even lightweight captioning and OCR tools can enhance rewriting quality. Furthermore, the 8-bit quantized version of this model delivers nearly equivalent performance to its full-precision counterpart—67.6% vs. 71.5% Intent Accuracy—while slightly improving Argument Similarity (79.5%), indicating its suitability for deployment on resource-constrained devices.
@@
Despite the strong performance of our ReVision variants, certain limitations hinder further accuracy gains. A primary issue is the loss of fine-grained text details caused by downsampling images to $256\times 256$ resolution, which impairs recognition of critical elements such as ingredient lists or nutritional facts on product packaging. Additionally, the dataset’s lack of explicit reference localization limits the model’s ability to align user intent with specific image regions, resulting in object disambiguation and instruction alignment errors. Future work could address these challenges by incorporating bounding box annotations to provide spatial grounding cues and by processing localized image regions rather than entire downsampled images, reducing information loss in text-heavy visual inputs. This approach aligns with Pali-Gemma’s short-resolution increase technique beyer2024paligemma, which improves fine-grained visual understanding. Nonetheless, our findings reaffirm that task-specific instruction tuning and metadata augmentation markedly enhance multimodal rewriting, supporting scalable and efficient on-device deployment.
@@ 7 Conclusion and Future Work
In this work, we explored Visual Instruction Rewriting as a lightweight, privacy-preserving approach to multimodal interaction on AR, VR, and smartphone devices. With a strong emphasis on dataset development, we present a diverse collection of 39,000+ examples covering 14 domains, enabling robust training for on-device instruction rewriting. Our approach ensures that text-only inference is more secure in privacy-sensitive settings by eliminating the need to send personal vision-related images to the server, reducing data exposure risks. Additionally, rewriting removes the necessity of storing images, making multimodal AI systems more efficient and privacy-focused. Our experimental results show that even an 8-bit quantized model maintains strong performance while significantly reducing memory requirements. For future work, we aim to expand data coverage by incorporating more diverse real-world multimodal instructions and introducing multilingual support to enhance accessibility. Furthermore, improving deixis resolution with bounding box annotations and localized image region training will enhance reference grounding, while integrating gaze tracking and tactile input can further refine contextual understanding in on-device AI assistants.
@@
While our approach demonstrates strong performance in Visual Instruction Rewriting, several limitations remain. First, image downsampling to $256\times 256$ resolution can lead to the loss of fine-grained text details, affecting instructions that rely on small-font information, such as nutritional labels or product specifications. Second, deictic reference resolution remains challenging, especially in images with multiple similar objects where the model lacks explicit localization cues. The absence of bounding box annotations in our dataset limits the model’s ability to disambiguate references, leading to errors in object-grounded instructions. Additionally, while our model is lightweight and optimized for on-device execution, it still lags behind larger VLMs in handling complex multimodal instructions requiring deep reasoning and external world knowledge. Lastly, our dataset, while diverse across 15+ domains, is monolingual, limiting applicability to multilingual and culturally varied settings. Future work can address these challenges by increasing dataset coverage, incorporating localized image region processing, and adding bounding box annotations to improve reference resolution and multimodal grounding.
Ethical Considerations